# Mixture of Directed Graphical Models for Discrete Spatial Random Fields

### v3

@@ 2.1 Compatibility of a graph to a NUG
In other words, we say a graph $\mathbf{G}$ is compatible with $\mathbf{N}$ when all the edges (directed or undirected) of $\mathbf{G}$ appear as undirected edges in $\mathbf{N}$. See Appendix A.2 for examples of compatible and incompatible DAGs for a given NUG. Note, that compatibility does not require that $|\mathbf{E}|=|\bar{\mathbf{E}}|$, where $|\mathbf{X}|$ denotes the cardinality of a set $\mathbf{X}$. We can extend this definition to sets of graphs. Let $\mathcal{G}=\{\mathbf{G}_{1},\dots,\mathbf{G}_{k}\}$ be a set of graphs. Then the set, $\mathcal{G}$, is compatible with $\mathbf{N}$, denoted $\mathcal{G}\Bumpeq\mathbf{N}$, if for each $\mathbf{G}\in\mathcal{G}$, $\mathbf{G}$ is compatible with $\mathbf{N}$.
The set of all graphs compatible with a NUG contains the set of all DAGs compatible with the NUG, which are of particular interest due to the simple and computationally efficient specification of a DGM given a DAG (see Section 3). We refer to complete set of DAGs compatible with a NUG as a super set, as it contains all the classes of DAGs (i.e. subsets of this super set) which we explore later in this Section. Let $\mathcal{D}(\mathbf{N})=\{\mathbf{D}:\mathbf{D}\Bumpeq\mathbf{N}\}$, be the set of all DAGs compatible with the NUG. A property of the association matrix for a DAG, $\mathbf{D}$, is that there exists a permutation of the rows of $A(\mathbf{D})$ for which $A(\mathbf{D})$ is lower triangular. Additionally, in a DAG, the $i$th row of $A(\mathbf{D})$ is a $m$-dimensional vector with binary indicators for the parents of $i$, that is $\pi(i)$.
2.2 Classes of Compatible DAGs
While the complete set of DAGs, $\mathcal{D}(\mathbf{N})$, compatible with the NUG comprises of the combined set all DAGs that can be derived from each possible undirected subgraph of the NUG, we define a class of compatible DAGs as a subset of the super set of compatible DAGs that are further distinguished by specific criteria. We explore two specific classes of compatible DAGs that are well studied in graph theory, namely acyclic orientations (AO) and spanning trees (ST). Let $\mathcal{D}^{\text{ST}}(\mathbf{N})$ denote the class of ST-DAGs and $\mathcal{D}^{\text{AO}}(\mathbf{N})$ denote the class of AO-DAGs. These two classes represent two ends of a spectrum of DAGs compatible with the NUG for which all vertices of the NUG remain connected in the DAGs of each class. By connected we mean that there exists a path between any two vertices in the graph. For the AO-DAGs, all edges of the NUG are preserved. In contrast, each of the ST-DAGs contain the minimal number of edges so that there is a path between any two vertices in $\mathbf{V}$. We also propose an intermediary class, which we call the rooted (R) class, denoted by $\mathcal{D}^{\text{R}}(\mathbf{N})$. The DAGs in the rooted class preserve most of the edges of the NUG and have a single orphan vertex from which all edges are oriented away. For all three of our classes, theory exists to count the cardinality of the class. The cardinality of the class will come into play in Section 3, when we define the MDGM prior. We review the algorithms for generating AO-DAGs and ST-DAGs from a NUG and present our algorithm of generating R-DAGs, all of which we utilize to obtain draws from the posterior distribution of our model.
2.2.1 Acyclic Orientations
compatible with the NUG, $\mathbf{N}=\{\mathbf{V},\mathbf{E}\}$, such that all undirected edges of the NUG appear as directed edges for every DAG in the set. A simple way to assign an acyclic orientation from the NUG is to permute the rows and columns of $A(\mathbf{N})$ and then take the lower triangular matrix as the association matrix for a DAG. As noted above, the number of acyclic orientations of an undirected graph is fewer than number of permutations of the vertices, as multiple permutations will correspond to the same acyclic orientation. While the theory exists to count and enumerate the number of acyclic orientations (Stanley, 1973; Squire, 1998), in practice both tasks are very computationally expensive, even for orderly graph structures such as a first-order regular lattice. For computational expediency, we will generate random acyclic orientations by permuting the indices in our MCMC algorithm in Section 4. We refer the reader to Appendix B.2 for the relevant theory to count the number of acyclic orientations of a graph.
2.2.2 Rooted Graphs
Begin with $i$ as the root vertex.
for $j\in\mathbf{V}$ do for $j\in\mathbf{V}$ do
      for $j\in\mathbf{V}$ do $l(j)\leftarrow min_{(i-j)}\sum_{\{s,r\}\in(i-j)}w(\{s,r\})$ $l(j)\leftarrow min_{(i-j)}\sum_{\{s,r\}\in(i-j)}w(\{s,r\})$
      $l(j)\leftarrow min_{(i-j)}\sum_{\{s,r\}\in(i-j)}w(\{s,r\})$triangleright  where $(i-j)$ denotes a path from vertex $i$ to vertex $j$.   $\triangleleft$ where $(i-j)$ denotes a path from vertex $i$ to vertex $j$.   $\triangleleft$ where $(i-j)$ denotes a path from vertex $i$ to vertex $j$.   $\triangleleft$
Define $\vec{\mathbf{E}}$ as:
for $\{r,s\}\in\bar{\mathbf{E}}$ do for $\{r,s\}\in\bar{\mathbf{E}}$ do
      for $\{r,s\}\in\bar{\mathbf{E}}$ do if $l(r)<l(s)$ then $(s,r)\in\vec{\mathbf{E}}$ if $l(r)<l(s)$ then $(s,r)\in\vec{\mathbf{E}}$
      if $l(r)<l(s)$ then $(s,r)\in\vec{\mathbf{E}}$ else if $l(r)>l(s)$ then $(r,s)\in\vec{\mathbf{E}}$ else if $l(r)>l(s)$ then $(r,s)\in\vec{\mathbf{E}}$
      else if $l(r)>l(s)$ then $(r,s)\in\vec{\mathbf{E}}$ else if $l(r)=l(s)$ then $\{r,s\}\not\in\vec{\mathbf{E}}$ else if $l(r)=l(s)$ then $\{r,s\}\not\in\vec{\mathbf{E}}$ else if $l(r)=l(s)$ then $\{r,s\}\not\in\vec{\mathbf{E}}$
The following describes Algorithm 1 to create an R-DAG for any NUG. First, assign weights to the edges of the NUG, $w(\{i,j\})$ for $\{i,j\}\in\bar{\mathbf{E}}$, as follows: edges that represent areal units with borders touching at more than a point assign the weight one, for areal units connected by borders touching at just a point assign the weight two. Then select a vertex, $i$, to be the root. Label all vertices by the smallest weighted path from the root vertex. Orient edges from lower-valued labels to higher-valued labels. Edges between vertices with the same label are deleted. The algorithm provides a class of compatible DAGs with a reduced number directed edges (therefor an intermediate between the AO and ST classes) and cardinality equal to the number vertices in the NUG.
2.2.3 Spanning Trees
Lastly, we introduce the class directed spanning trees. We limit $\mathcal{D}^{\text{ST}}(\mathbf{N})$ to be the set of all possible rooted and directed spanning trees (i.e. a directed spanning tree with a single root, referred to as a rooted tree from here out) generated from a NUG for reasons explained in Section 3. To generate a rooted tree from a NUG, first an undirected tree is generated using the loop erased random walk algorithm, which generates uniform draws of spanning trees from a NUG (Wilson, 1996). A loop erased random walk proceeds by performing a random walk along the edges of a graph to generate a path of vertices that does not contain any cycles (loops). If the random walk returns to a vertex that has already been visited in the path, then the cycle that is created by returning to vertex, $j$, is erased from the path and the random walk then proceeds from $j$. To obtain a spanning tree from a NUG, randomly select a vertex of the graph to be the first vertex of the tree. Then select another vertex at random and from this vertex perform a loop erased walk until the path of the random walk reaches the first vertex of the tree. This path and the vertices thereof are now part of the tree. Continue to select vertices at random that are not already part of the tree and perform a loop erased random walk until the path connects to vertices that are part of the tree. This algorithm guarantees a uniform draw of all possible spanning trees from a NUG (Wilson, 1996). We modify this algorithm in a novel MCMC scheme obtain draws of trees from the posterior. A rooted tree is then generated by selecting a vertex to be the root, and then orienting all edges away from the root.
Kirchhoff’s theorem allows us enumerate the the number of spanning trees that can be generated from an NUG through submatrices of the Laplacian matrix for a graph. The Laplacian of a graph is the degree matrix minus the adjacency matrix, $W(\mathbf{N})-A(\mathbf{N})$, where $W(\mathbf{N})=\mbox{diag}\left(A(\mathbf{N})\mathbf{1}_{n}\right)$ denotes the degree matrix. By Kirchhoff’s theorem, the number of spanning trees from a NUG, $\mathbf{N}$, is the cofactor of any $(i,j)$ position in the matrix. A $(i,j)$ cofactor is calculated by deleting the $i$th row and $j$th column of the Laplacian and taking the determinant times $(-1)^{i+j}$.
The number of ST-DAGs from any spanning tree is the number of vertices in the graph, thus the total number of ST-DAGs from the NUG is the number of spanning trees times the number of vertices; however, for reasons explained in Section 3 the effective size of the ST-DAG class reduces to simply the number of undirected spanning trees.
@@ 3 Hidden Discrete-Valued Spatial Dependence
where $I(\cdot)$ is the indicator function and the identifiability constraint, $\eta_{1}>\eta_{0}$, avoids the label switching problem of the latent variable. In the case of the of the MDGM, $\boldsymbol{\theta}=\{\mathbf{D},\boldsymbol{\xi}\}$, whereas $\boldsymbol{\theta}=\boldsymbol{\xi}$ for the MRF and aMRF. Let $p(\boldsymbol{\theta})=p(\mathbf{D})p(\boldsymbol{\xi})$ for the MDGM and $p(\boldsymbol{\theta})=p(\boldsymbol{\xi})$ for the models with MRF and aMRF priors. We use $\boldsymbol{\xi}$ to denote the spatial dependence parameters of the three different spatial models, and while the mathematical forms of the three latent variable priors that we use for analyses later in the paper are analogous, the mathematical interpretation of $\boldsymbol{\xi}$ is distinct between the MRF/aMRF and MDGM priors and across our highlighted classes of the MDGM. A default choice for the prior $\mathbf{D}$, is a uniform prior, $p(\mathbf{D})=\frac{1}{|\mathcal{D}^{\dagger}(\mathbf{N})|}$ for $\mathbf{D}\in\mathcal{D}^{\dagger}(\mathbf{N})$, where $\mathcal{D}^{\dagger}(\mathbf{N})\subseteq\mathcal{D}(\mathbf{N})$ is a specific class/subset of all compatible DAGs.
@@ 3.1 MDGM Prior
In the analyses which follow, for the ST and rooted MDGM priors, we specify the prior on the DAGs, $p(\mathbf{D})=\frac{1}{|\mathcal{D}^{\dagger}(\mathbf{N})|}$, to be uniform over all possible DAGs in the set $\mathcal{D}^{\dagger}(\mathbf{N})$. For computational simplicity in the AO case, we let the uniform draws of permutations induce a prior over acyclic orientations; that is, the prior over acyclic orientations is proportional to the number of permutations that correspond to each acyclic orientation, which is roughly uniform.
Additionally, for all three highlighted classes of the MDGM, we set
@@
$p(z_{i}|\mathbf{z}_{-i},\beta)=\frac{\exp\left(\beta\sum_{j\in\partial(i)}\mathbf{I}(z_{i}=z_{j})\right)}{\prod_{k\in\kappa(i)}\left[\exp\left(\beta\sum_{j\in\pi(k)}\mathbf{I}(z_{j}=0)\right)+\exp\left(\beta\sum_{j\in\pi(k)}\mathbf{I}(z_{j}=1)\right)\right]},$
@@
be the set of all children of vertex $i$. From the prior full conditionals we see that $p(z_{i}|\mathbf{z}_{-i})$ depends only the set of variables associated with the vertices $\{\pi(i),\kappa(i),\pi(\kappa(i))\}$, that is, the set of parents of $i$, children of $i$, and the parents of the children of $i$. This set is called the Markov blanket of $i$/$z_{i}$, or the minimal set of vertices/variables, that when conditioned on, make $i$/$z_{i}$ independent of all other vertices/variables. In the case of the rooted class, the set $\mathcal{D}^{R}(\mathbf{N})$ are Markov blanket equivalent with the NUG, $\mathbf{N}$, when the NUG is a regular lattice with a second-order neighborhood structure (see Appendix C.1 for more details). We define Markov blanket equivalence of a class such that for each $\mathbf{D}\in\mathcal{D}^{\dagger}(\mathbf{N})$ every vertex $i$ in the DAG has the same Markov blanket as the corresponding vertex $i$ in the NUG.
We also note a distinct property of the spanning trees in regards to identifiability of the DAG topology as alluded to in Section 2.2.3. It can be shown that specifying a probability distribution for a tree graph can be written equivalently as an MRF or as a DGM, where the DGM can be represented by a rooted tree (Meila and Jordan, 2001). Since the DGMs corresponding to the $n$ different rooted trees which can be created from an undirected tree are equivalent, the root of the rooted tree is unidentifiable, thus we are only able to learn the posterior distribution of the undirected tree structures. As a result, this class is less strictly a subset of DAGs when compared to the other two classes of DAG models; however, as a DGM and MRF specification for the same tree structure are equivalent, we can still utilize Equations 3 and 4 for the ST class to maintain unified computation between the MDGM classes.
@@
Lastly, we note the posterior, $p(\mathbf{z},\mathbf{D},\boldsymbol{\xi},\boldsymbol{\eta}|\mathbf{y})$, for a general MDGM and our highlighted classes as defined by Equation 2, is a valid probability distribution as each $p$ on the right hand side is a valid probability distribution. The particular form of the prior $p(\mathbf{z}|\mathbf{D},\boldsymbol{\xi})$ provides quick evaluation, needed for updates on $\mathbf{D}$ and $\boldsymbol{\xi}$ in the MCMC algorithms described in Section 4. In the case of the ST class, the only identifiable learning of the graph topologies is $\mathbf{U}(\mathbf{D})=\{\mathbf{V},\{\{i,j\}:(i,j)\in\vec{\mathbf{E}}\}\}$, the undirected structure of the DAGs in the posterior.

### v4

@@ 2.1 Weak compatibility of a graph to a NUG
See Appendix A.2 for examples of weakly compatible and incompatible DAGs for a given NUG. Note, that compatibility does not require that $|\mathbf{E}|=|\bar{\mathbf{E}}|$, where $|\mathbf{X}|$ denotes the cardinality of a set $\mathbf{X}$. We can extend this definition to sets of graphs. Let $\mathcal{G}=\{\mathbf{G}_{1},\dots,\mathbf{G}_{k}\}$ be a set of graphs. Then the set, $\mathcal{G}$, is weakly compatible with $\mathbf{N}$, denoted $\mathcal{G}\bumpeq\mathbf{N}$, if for each $\mathbf{G}\in\mathcal{G}$, $\mathbf{G}$ is weakly compatible with $\mathbf{N}$.
In the following section, we introduce strong compatibility and show that strongly compatible graphs are a subset of weakly compatible graphs. An alternative definition of weak compatibility is useful to this end and relies on the notion of a skeleton graph. The skeleton of a graph $\mathbf{G}$, denoted $\mathbf{U}(\mathbf{G})=\{\mathbf{V},\{\{i,j\}:(i,j)\in\mathbf{E}\text{ or }\{i,j\}\in\mathbf{E}\}\}$, is the undirected graph for which all directed edges of $\mathbf{G}$ become undirected. The following lemma provides an equivalent definition of weak compatibility.
Lemma 2.1.
A graph $\mathbf{G}$ is weakly compatible with $\mathbf{N}$ if and only if the skeleton of $\mathbf{G}$, $\mathbf{U}(\mathbf{G})$, is a subgraph of $\mathbf{N}$.
From Definition 2.1 and Lemma 2.1, graphs weakly compatible with a NUG cannot have edges between vertices that are not neighbors in the NUG. Strong compatibility tightens the structural criterion of weak compatibility further by considering conditional dependence relationships induced by directed edges.
2.2 Strong compatibility of a graph to a NUG
The appeal of graphical models is their representation of the dependence relationships between sets of variables in a probability distribution through a graph. In Appendix G we review how DAGs and undirected graphs can each express different sets of conditional independence relationships of the form $\mathbf{z}_{A}\perp\mathbf{z}_{B}|\mathbf{z}_{C}$ for sets of variables indexed by the sets vertices, $A,B,C\in\mathbf{V}$, in the vertex set. Let $I(\mathbf{G})$ denote the set of conditional independence relationships of the form $A\perp B|C$ asserted by the graph $\mathbf{G}$. To build up to our definition of strong compatibility which relates the sets $I(\mathbf{N})$ and $I(\mathbf{D})$ to each other, we first provide some common terms describing relationships between the sets of conditional independence relationships asserted by a graph and a probability distribution. Let $I(p)$ be the set of conditional independence relationships of the form, $A\perp B|C$, that hold in the probability distribution $p$, that is, $\mathbf{z}_{A}\perp\mathbf{z}_{B}|\mathbf{z}_{C}$. We say that $\mathbf{G}$ is an independence map (I-map) of $p$ if $I(\mathbf{G})\subseteq I(p)$; expressly, $\mathbf{G}$ does not assert any conditional independence relationships that are not present in the probability distribution (Pearl, 1988). When $I(p)\subseteq I(\mathbf{G})$, then we say that $\mathbf{G}$ is a dependence map (D-map) of $p$, accordingly, $\mathbf{G}$ does not assert any conditional dependencies that are not present in the probability distribution. When $I(\mathbf{G})=I(p)$, then $\mathbf{G}$ is a perfect map (P-map) of $p$.
We extend these definitions to relate two graphs to each other. For $\mathbf{G}$ and $\mathbf{G}^{\prime}$ defined on the same set of vertices, we say that $\mathbf{G}$ is a I-map of $\mathbf{G}^{\prime}$ or, equivalently, $\mathbf{G}^{\prime}$ is an D-map of $\mathbf{G}$ when $I(\mathbf{G})\subseteq I(\mathbf{G}^{\prime})$. When $I(\mathbf{G})=I(\mathbf{G}^{\prime})$, then $\mathbf{G}$ and $\mathbf{G}^{\prime}$ are said to be Markov equivalent (Andersson et al., 1997). The definition of strong compatibility is as follows.
Definition 2.2.
A graph, $\mathbf{G}=\{\mathbf{V},\mathbf{E}\}$, with vertices, $\mathbf{V}$, is strongly compatible with a NUG, $\mathbf{N}=\{\mathbf{V},\bar{\mathbf{E}}\}$, defined on the same set of vertices, denoted $\mathbf{G}\Bumpeq\mathbf{N}$, when $I(\mathbf{N})\subseteq I(\mathbf{G})$. That is, $\mathbf{G}$ is a D-map of $\mathbf{N}$.
As with weak compatibility, we say that a set of graphs, $\mathcal{G}$, is strongly compatible with $\mathbf{N}$, denoted $\mathcal{G}\Bumpeq\mathbf{N}$, when for each $\mathbf{G}\in\mathcal{G}$, we have $\mathbf{G}\Bumpeq\mathbf{N}$. The strong compatibility criterion ensures that compatible graphs do not introduce any dependencies between vertices that are not already present in the NUG. As a consequence, we have the following theorem with the proof in Appendix B.1.
Theorem 2.1.
Any graph $\mathbf{G}$ that is strongly compatible with a NUG $\mathbf{N}$, $\mathbf{G}\Bumpeq\mathbf{N}$, is also weakly compatible with $\mathbf{N}$, $\mathbf{G}\bumpeq\mathbf{N}$.
A DAG is weakly compatible as long as all of its directed edges correspond to undirected edges in the NUG. The complete set of DAGs weakly compatible with the NUG, $\mathcal{D}(\mathbf{N})=\{\mathbf{D}:\mathbf{D}\bumpeq\mathbf{N}\}$, comprises of the combined set of all DAGs that can be derived from each possible undirected subgraph of the NUG. Strong compatibility narrows the set of weakly compatible DAGs that simply honor the structure of the NUG to the set of DAGs which also honor the set of conditional independencies implied by the NUG. Conditional dependencies not present in the NUG in weakly compatible DAGs are introduced by v-structures. A v-structure is a pattern in a graph where two directed edges meet head-to-head, as in $i\rightarrow j\leftarrow k$. The parents of $j$, vertices $i$ and $k$, are called unprotected or unwed if there is no edge, directed or undirected, between $i$ and $k$. For such a pattern we have $i\perp k$, but $i\not\perp k|j$. Another characterization of strong compatibility is a weakly compatible DAG that also does not contain any v-structures with unwed parents. Trivially, the graph without any edges is strongly compatible with any NUG over the same set of vertices, though in the spatial setting such a graph is useless in learning and accounting for spatial dependence.
Since the sets of all weakly or all strongly compatible graphs are both quite large, we define a class of compatible DAGs as a subset of the complete set of compatible DAGs that are further distinguished by specific criteria. One criterion is DAGs that preserve all of the edges in the NUG. Such DAGs are called acyclic orientations with respect to the NUG and are a class of weakly compatible DAGs. We expound acyclic orientations and an additional class of DAGs in Appendix D. In the next section, we explore the class of spanning trees.
2.3 Spanning Trees
Spanning trees are fully connected graphs (i.e. graphs with a path between any two vertices) with a minimal edge set. In order to be fully connected, spanning trees have $n-1$ edges. Spanning trees are a unique class of compatible graphs: both undirected spanning trees (or simply, spanning trees) and a special case of directed spanning trees represent equivalent sets of conditional independence relationships. A directed spanning tree is a DAG with a spanning tree skeleton. Rooted spanning trees are a special case of directed spanning trees that have a single orphan vertex which is called the root of the graph (a generic DAG is also called rooted if it has a single orphan vertex). A rooted spanning tree does not contain any v-structures because all edges are oriented away from the root giving each non-root vertex only a single parent. Let $\mathcal{D}^{\text{ST}}(\mathbf{N})$ be the set of all possible rooted spanning trees.
Any two graphs that have the same unwed v-structures and skeleton graph are Markov equivalent (Verma and Pearl, 1990). Accordingly, a spanning tree is Markov equivalent with any rooted spanning tree that has the same skeleton. The Markov equivalence of a spanning tree and any rooted spanning tree with an identical skeleton has implications for statistical learning which we explain in Section 3. Due to the Markov equivalence, we call the class simply spanning trees, but leverage rooted spanning trees for unified computation in and presentation of the MDGM framework.
Spanning trees are also a special class of compatible graphs, which we formalize in the following theorem.
Theorem 2.2.
A spanning tree, $\mathbf{U}=\{\mathbf{V},\bar{\mathbf{E}}^{\prime}\}$, derived from a NUG, $\mathbf{N}=\{\mathbf{V},\bar{\mathbf{E}}\}$, is always strongly compatible with $\mathbf{N}$.
The proof is provided in Appendix B.2. Rooted spanning trees are also strongly compatible with the NUG because of the Markov equivalence of rooted and undirected spanning trees. That is, for all $\mathbf{D}\in\mathcal{D}^{\text{ST}}(\mathbf{N})$ we have $I(\mathbf{N})\subseteq I(\mathbf{D})=I(\mathbf{U}(\mathbf{D}))$. Rooted spanning trees honor the dependence relationships of the NUG because they do not contain unprotected v-structures. This naturally leads to the following corollary of Theorem 2.2.
Corollary 2.1.
For a NUG, $\mathbf{N}$, with first-order dependence structure on a regular lattice, a spanning tree is the only fully connected graph structure that is strongly compatible with $\mathbf{N}$.
The proof is in Appendix B.3. The strong compatibility of spanning trees makes the class a good default option for specifying an MDGM model. We explain how to specify a MDGM model given a general set of DAGs in Section 3.1 and highlight the spanning tree class. In Appendix C, we describe how to generate a rooted spanning tree and how to count the size of class.
@@ 3 Hidden Discrete-Valued Spatial Dependence
where $I(\cdot)$ is the indicator function and the identifiability constraint, $\eta_{1}>\eta_{0}$, avoids the label switching problem of the latent variable. In the case of the MDGM, $\boldsymbol{\theta}=\{\mathbf{D},\boldsymbol{\xi}\}$, whereas $\boldsymbol{\theta}=\boldsymbol{\xi}$ for the MRF and aMRF. Let $p(\boldsymbol{\theta})=p(\mathbf{D})p(\boldsymbol{\xi})$ for the MDGM and $p(\boldsymbol{\theta})=p(\boldsymbol{\xi})$ for the models with MRF and aMRF priors. We use $\boldsymbol{\xi}$ to denote the spatial dependence parameters of the three different spatial models, and while the mathematical forms of the three latent variable priors that we use for analyses later in the paper are analogous, the mathematical interpretation of $\boldsymbol{\xi}$ is distinct between the MRF/aMRF and MDGM priors and across our classes of the MDGM highlighted in the paper and appendices. In the MRF/aMRF prior, spatial dependence between the vertices of a NUG is determined solely through the parameters $\boldsymbol{\xi}$, whereas in the MDGM classes, spatial dependence is determined through the DAGs included in the mixture and the parameters $\boldsymbol{\xi}$. A default choice for the prior $\mathbf{D}$, is a uniform prior, $p(\mathbf{D})=\frac{1}{|\mathcal{D}^{\dagger}(\mathbf{N})|}$ for $\mathbf{D}\in\mathcal{D}^{\dagger}(\mathbf{N})$, where $\mathcal{D}^{\dagger}(\mathbf{N})\subseteq\mathcal{D}(\mathbf{N})$ is a specific class/subset of all weakly compatible DAGs.
@@ 3.1 MDGM Prior
Additionally, for a general DAG we set
@@
$p(z_{i}|\mathbf{z}_{-i},\beta)\propto\frac{\exp\left(\beta\sum_{j\in\partial(i)}\mathbf{I}(z_{i}=z_{j})\right)}{\prod_{k\in\kappa(i)}\left[\exp\left(\beta\sum_{j\in\pi(k)}\mathbf{I}(z_{j}=0)\right)+\exp\left(\beta\sum_{j\in\pi(k)}\mathbf{I}(z_{j}=1)\right)\right]},$ (5)
@@
be the set of all children of vertex $i$. From the prior full conditionals we see that $p(z_{i}|\mathbf{z}_{-i})$ depends only the set of variables associated with the vertices $\{\pi(i),\kappa(i),\pi(\kappa(i))\}$, that is, the set of parents of $i$, children of $i$, and the parents of the children of $i$.
@@
We note that the posterior, $p(\mathbf{z},\mathbf{D},\boldsymbol{\xi},\boldsymbol{\eta}|\mathbf{y})$, for a general MDGM as defined by Equation 2, is a valid probability distribution as each $p$ on the right hand side is a valid probability distribution. The particular form of the prior $p(\mathbf{z}|\mathbf{D},\boldsymbol{\xi})$ provides quick evaluation, needed for updates on $\mathbf{D}$ and $\boldsymbol{\xi}$ in the MCMC algorithms described in Section 4. We now highlight the spanning tree class.
3.1.1 Spanning Tree Class
First, we specify a prior for the spanning tree class. As noted above when describing the loop erased random walk algorithm, the probability of a spanning tree is proportional to the product of the weights of its edges. Specifying uniform weights for the edges of the NUG, a priori, yields the uniform prior $p(\mathbf{D})=\frac{1}{|\mathcal{D}^{\text{ST}}(\mathbf{N})|}$, over all possible spanning trees in the set $\mathcal{D}^{\text{ST}}(\mathbf{N})$. Prior information on edge inclusion probabilities can be included into the model by assigning different weights to the edges in the NUG. Also of note, in the prior full conditionals for the spanning tree class, the denominator of Equation 5 becomes a constant, $2(1+\exp(\beta))^{n-1}$, because a rooted spanning tree does not have any unprotected v-structures.
Consider now the characterization of spatial dependence in the spanning tree class. The pairwise spatial dependence between two vertices is determined by the proportion of times the two vertices are connected across all spanning trees in the mixture and the spatial dependence parameter $\beta$. We can think of $\beta$ as the global spatial dependence parameter and the weight edges as local adjustments to spatial dependence. As $\beta$ increases, the areal units will be more clustered. The probability that a particular edge appears in any given spanning tree in the mixture is proportional to the weight of that edge. With uniform prior weights on the edges in the spanning tree class, A priori, all edges have equal probability of appearing in the mixture, thus there are no local adjustments to spatial dependence. For nonuniform edge weights, as in the posterior distribution, local adjustments to the degree of dependence do occur, making the spanning tree class very flexible in representing spatial dependence.
Lastly, we address the previously alluded to property of the spanning tree class in regard to identifiability of the DAG topology. It can be shown that specifying a probability distribution for a tree graph can be written equivalently as an MRF or as a DGM, where the DGM is represented by a rooted spanning tree (Meila and Jordan, 2001). Since the DGMs corresponding to the $n$ different rooted spanning trees which can be created from an undirected tree are Markov equivalent, the root of the rooted spanning tree is unidentifiable, thus we are only able to learn the posterior distribution of the undirected tree structures. As a result, this class is less strictly a subset of DAGs when compared to the other two classes of DAG models that we present in the appendix; however, as a DGM and MRF specification for the same tree structure are equivalent, we can still utilize Equations 3 and 4 for the spanning tree class to maintain unified computation between the MDGM classes.